# Knowledge Prompting: How Knowledge Engineers Use Large Language Models

### v1

@@ Abstract.
Despite many advances in knowledge engineering (KE), challenges remain in areas such as engineering knowledge graphs (KGs) at scale, keeping up with evolving domain knowledge, multilingualism, and multimodality. Recently, KE has used LLMs to support semi-automatic tasks, but the most effective use of LLMs to support knowledge engineers across the KE activites is still in its infancy. To explore the vision of LLM copilots for KE and change existing KE practices, we conducted a multimethod study during a KE hackathon. We investigated participants’ views on the use of LLMs, the challenges they face, the skills they may need to integrate LLMs into their practices, and how they use LLMs responsibly. We found participants felt LLMs could contribute to improving efficiency when engineering KGs, but presented increased challenges around the already complex issues of evaluating the KE tasks. We discovered prompting to be a useful but undervalued skill for knowledge engineers working with LLMs, and note that natural language processing skills may become more relevant across more roles in KG construction. Integrating LLMs into KE tasks needs to be mindful of potential risks and harms related to responsible AI. Given the limited ethical training, most knowledge engineers receive solutions such as our suggested ‘KG cards’ based on data cards could be a useful guide for KG construction. Our findings can support designers of KE AI copilots, KE researchers, and practitioners using advanced AI to develop trustworthy applications, propose new methodologies for KE and operate new technologies responsibly.
Knowledge Graph, Knowledge Engineering, Large Language Models, Hackathon, Ethnographic Study, Interviews, Knowledge Engineers Skills, Bias
$c$s: Computing methodologies Ontology engineeringccs: Human-centered computing Natural language interfacesccs: Human-centered computing Ethnographic studies
1. Introduction
Knowledge engineering (KE) is the process of capturing, structuring, representing and maintaining knowledge in a machine-readable way (Studer et al., 1998). The outcome of this process often takes the form of a knowledge graph (KG) (Hogan et al., 2022). Despite the remarkable evolution of the field, some challenges remain. In particular, that of working at scale across all tasks, from extracting structured knowledge from large collections of domain information to keeping up with how domain knowledge evolves, to supporting multiple languages and modalities to assessing quality. The solution to scaling lies in automation, for which large language models (LLMs) hold great promise (Noy et al., 2019; Peng et al., 2023).
LLMs have changed the way we work in many fields, including KE. Knowledge engineers have been using natural language processing (NLP) methods and tools to semi-automate tasks like information extraction. However, we don’t know how LLMs affect or could affect those practices. To explore the changes in KE practices, we studied the interaction of researchers with LLMs to solve KE tasks. For this purpose, the idea of a hackathon emerged during the Dagstuhl Seminar on “Knowledge Graphs and their Role in the Knowledge Engineering of the 21st Century” held in September 2022.111https://www.dagstuhl.de/22372 We invited researchers and practitioners from the AI and KE fields to a four-day hackathon to investigate LLMs’ automation support for KE tasks. 222urlhttps://king-s-knowledge-graph-lab.github.io/knowledge-prompting-hackathon/
Our study explored what challenges knowledge engineers face, the skills they may need to have, and what safety testing means in this concept. Our investigation included an ethnographic study during the hackathon, followed by a set of semi-structured interviews after the event and a review of the documentary output of the hackathon. Our research questions are:
  1. (1)
@@ 1. Introduction
    What are the main challenges experienced by knowledge engineers when using LLMs for KE tasks?
  2. (2)
    How do knowledge engineers evaluate LLMs’ output for their practices?
  3. (3)
@@
    What skills does a knowledge engineer need to incorporate LLMs into their practice?
  4. (4)
    How aware are knowledge engineers of using LLMs responsibly?
  5. (5)

### v3

@@ Abstract
Despite many advances in knowledge engineering (KE), challenges remain in areas such as engineering knowledge graphs (KGs) at scale, automating tasks, and keeping pace with evolving domain knowledge. KE has used NLP demonstrating notable advantages in knowledge-intensive tasks, but the most effective use of generative AI to support knowledge engineers across the KE activities is still in its infancy. To explore how generative AI may enhance KE and change existing KE practices, we conducted a multi-method study during a KE hackathon. We investigated participants’ views on the use of generative AI, the challenges they face, the skills they may need to integrate generative AI into their practices, and how they use generative AI responsibly. We found participants felt LLMs could indeed contribute to improving efficiency when engineering KGs, but presented increased challenges around the already complex issues of evaluating KE task success. We discovered prompting to be a useful but undervalued skill for knowledge engineers working with LLMs, and note that NLP skills may become more relevant across more roles in KE workflows. Integrating generative AI into KE tasks needs to be done with awareness of potential risks and harms. Given the limited ethical training most knowledge engineers receive, solutions such as our proposed ‘KG Cards’ based on Data Cards could be a useful guide for KG construction. Our findings can support designers of KE AI copilots, KE researchers, and practitioners using advanced AI to develop trustworthy applications, propose new methodologies for KE and operate new technologies responsibly.
keywords:
[…]
{highlights}
Large Language Models (LLMs) can enhance efficiency in knowledge engineering tasks like knowledge graph construction, but their effective use requires new skills, such as advanced prompting and responsible AI integration.
The introduction of tools such as KG Cards can address ethical and transparency challenges in LLM-assisted knowledge engineering by guiding practitioners towards safer and more responsible practices.
Generative AI offers promising potential for assisting less experienced users in KE tasks, but more research is needed to mitigate biases and ensure reliability throughout the KG development lifecycle.
1 Introduction
Knowledge engineering (KE) is the process of capturing, representing and maintaining knowledge in a machine-readable way. It involves the development of knowledge-based systems (KBSs), and tools and principles to operationalise these [163, 32]. Since Google launched its knowledge graph (KG) over a decade ago, KGs (especially large ones with millions of entities and billions of facts in triples) have become the predominant way to represent, store, and utilise knowledge, offering several advantages across various domains. Recebtly they have been used to enhance pre-trained language models (PLMs) [100] and retrieval-augmented generation (RAG) systems [90] and offer explanations in artificial intelligence (AI) applications [167]. The techniques used for engineering these KGs have benefited from development in Natural Language Processing (NLP) techniques across tasks such as knowledge extraction, completion, and inconsistency detection [149]. More recently, the most advanced branch of NLP, large language models (LLMs), such as GPT models [24, 121] and the LLaMA series [168, 169], have demonstrated notable potential for knowledge-intensive tasks, with applications for KG construction through methods such as knowledge probing [132, 6, 193]. The integration of KGs and LLMs has become a growing trend that fosters a mutually beneficial relationship: LLMs enhance KG construction and maintenance, while KGs are used to train, prompt, augment, and evaluate LLMs [123].
Although research efforts to unify LLM and KG have gained popularity, debates and concerns continue. Critical perspectives around misinformation, ethical issues, privacy and security, bias, transparency, accountability and many others have emerged at a broader level and apply to LLM-based systems [124, 128]. Regardless of improvements in automated processes, it may be that KE will always require some form of human-in-the-loop participation to build trust with end-users and stakeholders [58].
Despite the increasing adoption of LLMs in KE, there is a notable gap in user studies focusing on the user-centric aspects of LLM technology in KE.To explore the opportunities and transformations in KE practices enabled by generative AI, it is essential to analyse how researchers and knowledge engineers interact with this emerging technology in standard KE tasks. To this end, the idea of a hackathon focused on LLMs in KE originated during the Dagstuhl Seminar on “Knowledge Graphs and their role in the Knowledge Engineering of the 21st Century” held in September 2022 [58]. [^1] https://www.dagstuhl.de/22372. Building on this idea, we organised a four-day hackathon, inviting researchers and practitioners from the AI and KE communities to explore generative AI’s potential in automating KE tasks. [^2] https://king-s-knowledge-graph-lab.github.io/knowledge-prompting-hackathon/
The user study presented in this work leverages the hackathon as a test bed. Our investigation included an ethnographic study conducted during the event, complemented by post-event semi-structured interviews and a review of the hackathon’s documentary outputs. During the four-day workshop, we observed knowledge engineers working on KE research tasks using generative AI models. KE is a broad field involving numerous tasks, tools and practices. In this study, we aim address a set of curated topics such as knowledge graphs and their construction with humanin-the-loop. Specifically, we examined the obstacles that participants encountered, the skills they needed, and the safety testing practices they used. Our research questions are:
  1. 1.
@@ 1 Introduction
    What are the main challenges experienced by knowledge engineers when using generative AI for KE tasks?
  2. 2.
    How do knowledge engineers evaluate generative AI output for their practices?
  3. 3.
@@
    What skills do knowledge engineers need to incorporate generative AI into their practice?
  4. 4.
    How aware are knowledge engineers of using generative AI responsibly?
  5. 5.